Minutes of the HERA Meeting 20.3.2006
HERA (B.Holzer):
Lumi production: about 6 pb^-1 in the last week, good duty cycle.
Main power station A failed Tue night (8.3.). By Rearranging supplies,
lumi operation was possible. Repaired on Wed, but the station will need
some major repair work in summer (down for 2 weeks, spare part ordered).
A proton quadrupole QC53 WL40 was caught with fast excursions, possibly
causing spikes. A fuse was exchanged, but this did not cure the problem
completely.
Pol tuning in progress: 1st order bumps done, 2nd order underway, energy
scan to be done.
ZEUS (U.Schneekloth):
Took data with varying efficiency due to beam conditions.
Integrated 8 pb^-8 for physics in 2006, out of 12 delivered.
e-collimator file adjusted last week (was too tight).
Some fills suffered from p background, possibly QC spikes.
H1 (J.Bracinik):
Took 4.3 pb^-1 with HV on, out of 6 delivered in 1 week.
2 fills on Monday had many proton spikes.
Roman pots are active, VFPS without extra p bump for increased
acceptance. VFPS frequently retracted due to high backgrounds.
Hermes (period coordinator):
Very good e beam condtions.
The solenoid (from now on called The Recoil Magnet) is ramped during
e- acceleration, with fast orbit stabilisator active at HERA.
Operation not yet stable, 2 quenches last week, He pump had to be
repaired, still need spare valve.
High density run operation has started, usually with hydrogen, which
makes background in H1 and ZEUS, tolerable at 13 mA.
Plan: routine lumi operation, increase e current towards 40 mA.
ZEUS is coordinating experiment.
Daniel Pitzl